T is an unbiased estimator for theta, show that
T^2 is a biased estimator for theta^2 .
1
Expert's answer
2020-09-07T18:06:34-0400

We have by assumption, E(T)="\theta"..........(Eq. 1)


Hence, "E(T^2)=Var(T)+(E(T))^2"


Substuting Eq. 1 in above equation we get,

"E(T^2)=Var(T)+\theta^2".................(Eq. 2)

"\therefore\ Var(T)=E(T^2)-\theta^2" ...........(Eq. 3)


But since Var(T)>0

"\therefore\ E(T^2)-\theta^2>0"

"\therefore E(T^2)>\theta^2"


Which means, "E(T^2) \not=\theta^2"


Hence we can say that "T^2" is a biased estimator of "\theta^2."